A British Thermal Unit (BTU) is the amount of heat required to raise
one (1) pound of water one (1) degree Fahrenheit. How many BTU’s
would be required to raise 45 pounds of water 18 degrees Fahrenheit?

Answer by CharlesG2(834)

A British Thermal Unit (BTU) is the amount of heat required to raise
one (1) pound of water one (1) degree Fahrenheit. How many BTU’s
would be required to raise 45 pounds of water 18 degrees Fahrenheit?
1 BTU/(1 lb * 1 F)=x BTU/(45 lb * 18 F)
1/1=x/(45*18)=x/810
1=x/810
x=810
810 BTU